MASS ANALYZED THRESHOLD IONIZATION OF LUTETIUM DIMER

\underline{LU WU}, MOURAD ROUDJANE, YANG LIU AND DONG-SHENG YANG, Department of Chemistry, University of Kentucky, Lexington, KY 40506-0055.

Lu_{2} is produced in a pulsed laser-vaporization metal-cluster source and studied by mass-analyzed threshold ionization (MATI) spectroscopy. The MATI spectrum displays several long progressions from the transitions between various vibrational levels of the neutral and ion electronic states. From the spectrum, the upper limit of the ionization energy of the dimer is determined to be 43996 cm^{−1}, and the vibrational frequencies are measured to be 121 cm^{−1} in the neutral state and 90 cm^{−1} in the ion state. By combining with ab initio calculations at CASPT2 level, the ground state of Lu_{2} is identified as ^{3}Σ_{g}^{−}. The ^{3}Σ_{g}^{−} state has an electron configuration of 6sσ_{g}^{2}5dπ_{u}^{1}5dπ_{u}^{1}6sσ_{u}^{2}, which is formed by the interactions of two Lu atoms in the ^{2}D(5d6s^{2}) ground state. Ionization of the neutral state removes a 5dπ_{u} bonding electron and yields a ion state with a considerably longer bond distance. Lu_{2} has a very different bonding feature from La_{2}, for which a ^{1}Σ_{g}^{+} ground state was previously identified with an electron configuration of 5dπ_{u}^{4}6sσ_{g}^{2} formed by the interactions of two La atoms in the ^{4}F(5d^{2}6s) excited state.^{a}

aYang Liu, Lu Wu, Chang-Hua Zhang, Serge A. Krasnokutski, and Dong-Sheng Yang, J. Chem. Phys. 135, 034309 (2011).